OpenStudy (anonymous):

really need help :( A sandbag was thrown downward from a building . The function f(t)=-16t^2-32+128 shows the height f(t), in feet of the sandbag after t seconds. Part A. Factor the function f(t) and use the factors to interpret the meaning of the x-intercept of the function. Part B. Complete the square of the expression for f(x) to determine the vertex of the graph of f(x) . Would this be a max or min what is the axis of symm

jimthompson5910 (jim_thompson5910):

I'll get you started on part A -16t^2-32t+128 -16*t^2-16*2t+16*8 -16(t^2+2t-8) ... factor out the GCF 16 and let you finish up

OpenStudy (anonymous):

is that how I would get a and b

jimthompson5910 (jim_thompson5910):

What do you get when you factor t^2+2t-8

OpenStudy (anonymous):

that's what im having trouble with I cant take out anything

jimthompson5910 (jim_thompson5910):

find two numbers that multiply to -8 (last term) and add to 2 (middle coefficient)

OpenStudy (anonymous):

I got (t-2) ( t+4)

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

therefore, -16t^2-32t+128 factors to -16(t-2)(t+4)

jimthompson5910 (jim_thompson5910):

how can you use this to find the x-intercepts?

OpenStudy (anonymous):

do I plug in o for t?

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

more like you set it equal to zero and solve for t

jimthompson5910 (jim_thompson5910):

-16t^2-32t+128 = 0 -16(t-2)(t+4) = 0 keep going...

OpenStudy (anonymous):

o ok thank you

jimthompson5910 (jim_thompson5910):

you're welcome

OpenStudy (anonymous):

is it 96t+32

jimthompson5910 (jim_thompson5910):

-16(t-2)(t+4) = 0 (t-2)(t+4) = 0 t-2 = 0 or t+4=0 solve each for t

OpenStudy (anonymous):

those are both the x intercepts>

jimthompson5910 (jim_thompson5910):

which are what?

OpenStudy (anonymous):

t= 2 and t=-4

jimthompson5910 (jim_thompson5910):

yep, the x-intercepts are 2 and -4 (basically the graph crosses the x axis at 2 and -4)

OpenStudy (anonymous):

your the best!!! can you help me with part b and c

jimthompson5910 (jim_thompson5910):

Part B. Complete the square of the expression for f(x) to determine the vertex of the graph of f(x) . Would this be a max or min

jimthompson5910 (jim_thompson5910):

To find the vertex, you first need to find the axis of symmetry

jimthompson5910 (jim_thompson5910):

how do we do this?

OpenStudy (anonymous):

I suck at math idk :(

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

hint: for ax^2 + bx + c, the axis of symmetry is x = -b/(2a)

OpenStudy (anonymous):

x=-1 ?

OpenStudy (anonymous):

sorry for asking all these questions.. its just all this is due tonight @jim_thompson5910

jimthompson5910 (jim_thompson5910):

think of -16t^2-32t+128 as -16x^2-32x+128 what are a, b, c?

OpenStudy (anonymous):

a is 16 b is -32 and c is 128

jimthompson5910 (jim_thompson5910):

a = -16 b = -32 c = 128

jimthompson5910 (jim_thompson5910):

x = -b/(2a) x = -(-32)/(2*(-16)) x = 32/(-32) x = -1 So you are correct. The axis of symmetry is x = -1

jimthompson5910 (jim_thompson5910):

Now plug this into y = -16x^2-32x+128 to find y

OpenStudy (anonymous):

is y =144

jimthompson5910 (jim_thompson5910):

yep

jimthompson5910 (jim_thompson5910):

for the vertex: the x coordinate is -1, the y coordinate is 144 so (h,k) = (-1,144) h = -1 k = 144

jimthompson5910 (jim_thompson5910):

a = -16 h = -1 k = 144 so... y = a(x-h)^2 + k y = -16(x-(-1))^2 + 144 y = -16(x+1)^2 + 144

jimthompson5910 (jim_thompson5910):

which means -16x^2-32x+128 transforms into y = -16(x+1)^2 + 144 after completing the square

OpenStudy (anonymous):

so what would I put for b

jimthompson5910 (jim_thompson5910):

what's the vertex?

OpenStudy (anonymous):

-1. 144

jimthompson5910 (jim_thompson5910):

(-1,144)

jimthompson5910 (jim_thompson5910):

The y coordinate of that vertex is 144 So that's the min or the max. Which one is it? Do you know how you can tell?

OpenStudy (anonymous):

its that max because it positive right?

jimthompson5910 (jim_thompson5910):

no, the y coordinate of the vertex doesn't determine min or max look at the value of 'a'

OpenStudy (anonymous):

its -16 so its a min?

jimthompson5910 (jim_thompson5910):

it's actually the opposite (yeah confusing I know) if a < 0, then we have a max if a > 0, then we have a min

OpenStudy (anonymous):

o so its a ma x haha

jimthompson5910 (jim_thompson5910):

therefore, y = 144 is the max

OpenStudy (anonymous):

thank you so much you don't know how much your helping me man

OpenStudy (anonymous):

I have one more set of questions if your not to busy

jimthompson5910 (jim_thompson5910):

Sure I can do one more question

OpenStudy (anonymous):

THE FUNCTION H(t)=-16t^2+vt+s shows the height H (t) in feet of a projectile launched vertically from s feet above the ground after y seconds. The intila speed of the projectile is v feet per second

OpenStudy (anonymous):

Part A. the projectile was launched from a height of 100 feet with an intial velocity of 60 feet per second. Create an equation to find the time taken by the projectile to fall on the ground.

OpenStudy (anonymous):

Part b. What is the maxium height that the projectile will reach?

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

@agent0smith

OpenStudy (anonymous):

plz someone help @agent0smith @jim_thompson5910

jimthompson5910 (jim_thompson5910):

For the equation H(t)=-16t^2+vt+s, v = initial launch velocity s = initial launch height

OpenStudy (anonymous):

thanks @jim_thompson5910

jimthompson5910 (jim_thompson5910):

so what is the equation?

OpenStudy (anonymous):

i, trying to do it but i don't know how to get it

jimthompson5910 (jim_thompson5910):

what are v and s?

OpenStudy (anonymous):

60 and s is 100

jimthompson5910 (jim_thompson5910):

so H(t)=-16t^2+vt+s turns into ????

OpenStudy (anonymous):

16(60)^2+160+100

jimthompson5910 (jim_thompson5910):

v = 60 s = 100

jimthompson5910 (jim_thompson5910):

we haven't said what the value of t was yet

OpenStudy (anonymous):

so 16t^2+60(t)+100

jimthompson5910 (jim_thompson5910):

yes, H(t)=-16t^2+vt+s becomes H(t)=-16t^2+60t+100

jimthompson5910 (jim_thompson5910):

now solve -16t^2+60t+100 = 0 to find when the projectile will hit the ground

OpenStudy (anonymous):

how would i solve that

jimthompson5910 (jim_thompson5910):

by using the quadratic formula

OpenStudy (anonymous):

so its 60 + or - square root of -60^2 - 4(-16) (100)/2(-16)

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

that -60^2 should be (60)^2

jimthompson5910 (jim_thompson5910):

but otherwise, so far, so good

OpenStudy (lena772):

How did you get 144 for -16*-1^2+32*-1+128??

OpenStudy (lena772):

I keep getting 112. @jim_thompson5910 @jayhawk31

jimthompson5910 (jim_thompson5910):

The original function is f(t)=-16t^2-32t+128 Plug t = -1 into this function to get f(t)=-16t^2-32t+128 f(-1)=-16(-1)^2-32(-1)+128 f(-1)=144 So you just forgot to place -1 in parenthesis and it's -32 (not +32)

OpenStudy (lena772):

oh okay thanks, i realized my mistake after :3

jimthompson5910 (jim_thompson5910):

That looks correct

OpenStudy (lena772):

Thank you

jimthompson5910 (jim_thompson5910):

np